OpenStudy (afloridagirl):

A punter kicks a football upward with an initial speed of 48 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the initial speed.

OpenStudy (afloridagirl):

is it three?

OpenStudy (afloridagirl):

@agent0smith

OpenStudy (welshfella):

when ball hits the ground h = 0

OpenStudy (afloridagirl):

so 0?

OpenStudy (welshfella):

solve 0 = 48t - 16t^2 t will be one root - corresponding to when the ball was kicked Its the other root you want.

OpenStudy (afloridagirl):

oh

OpenStudy (welshfella):

factoring t(48 - 16t) = 0

OpenStudy (welshfella):

so t = 0 and 48 - 16t = 0

OpenStudy (afloridagirl):

so its 0?

OpenStudy (welshfella):

solve the second equation and you have the answer

OpenStudy (afloridagirl):

3?

OpenStudy (welshfella):

No - you want value of t when ball hits the ground after trvalling through the air

OpenStudy (welshfella):

48 - 16t = 0 16t = 48 t = 3 yea

OpenStudy (afloridagirl):

so wait is it or isnt it three?

OpenStudy (afloridagirl):

im sorry im hella confused at this moment

OpenStudy (welshfella):

yes it is 3 the solution to 48-16t = 0 is t = 3

OpenStudy (afloridagirl):

okay thank you

OpenStudy (welshfella):

t = 0 seconds corresponds to when the ball is kicked and after 3 seconds it lands on ground again.

OpenStudy (welshfella):

- in both cases h = 0.

OpenStudy (afloridagirl):

so wait whats the answer?

OpenStudy (welshfella):

the question asks for the time taken to reach the ground after being kicked.

OpenStudy (afloridagirl):

and that is three?

OpenStudy (welshfella):

right

OpenStudy (afloridagirl):

okay thank you

OpenStudy (welshfella):

yw